Question 5 : The temperature remains constant as energy is added to a substance. How may the substance be changing?
From a gas to a solid
From a liquid to a solid
Change in the amoung of kinetic energy
From a liquid to a gas
Question 6 : The heat of combustion for 1 mole of carbon to carbon dioxide is -410 kJ. How many kJ of heat would be liberated from the complete combustion of 60.0 g of carbon?
-24,600 kJ
-82 kJ
-2050 kJ
650.2 kJ
1
Expert's answer
2017-03-04T07:19:05-0500
Question 5: The correct answer: from a liquid to a solid.
Question 6: C + O2 = CO2
n(C)=m(C)/M(C) =60/12=5 mol
If 1mol – (-410) kJ, that 5 mol – x kJ. So: x=(5∙(-410))/1=-2050 kJ.
So the correct answer: -2050 kJ
Answer: Question 1: from a liquid to a solid;
Question 2: -2050 kJ.